OpenStudy (anonymous):

I just don't understand. In the product F=qVxB , take q = 5, V= 2.0i + 4.0j + 6.0k and F = -170i + 70j + 10k. What then is in unit-vector notation if Bx = By?

OpenStudy (anonymous):

lets matrix play

OpenStudy (anonymous):

\[F = qV \times B = q\left[\begin{matrix}i & j & k\\ Vx & Vy & Vz\\Bx & By & Bz\end{matrix}\right]\] Plug in components of V as given by the equation, and q is a scalar multiple. The cross product of V and B can be given by the determinant of the matrix. Keeping variables for Bx, By, and Bz, use vector components to solve for Bx By and Bz separately,